### v2

# Emilia: An Extensive, Multilingual, and Diverse Speech Dataset for Large-Scale Speech Generation

Abstract
Recently, speech generation models have made significant progress by using large-scale training data. However, the research community struggle to produce highly spontaneous and human-like speech due to the lack of large-scale, diverse, and spontaneous speech data. This paper presents Emilia, the first multilingual speech generation dataset from in-the-wild speech data, and Emilia-Pipe, the first open-source preprocessing pipeline designed to transform in-the-wild speech data into high-quality training data with annotations for speech generation. Emilia starts with over 101k hours of speech in six languages and features diverse speech with varied speaking styles. To facilitate the scale-up of Emilia, the open-source pipeline Emilia-Pipe can process one hour of raw speech data ready for model training in a few mins, which enables the research community to collaborate on large-scale speech generation research. Experimental results validate the effectiveness of Emilia. Demos are available at: https://emilia-dataset.github.io/Emilia-Demo-Page/.
1 Introduction
In the past few years, the research of speech generation has made significant advancements with the emergence of various generative models and the use of large-scale training data. The models such as Vall-E[1], SoundStorm [2], VoiceBox [3] and NaturalSpeech 3 [4] have considerably progressed in (zero-shot) speech generation by considerably scaling up both the datasets and model sizes, achieving high similarity, voice quality, and naturalness on academic datasets [5]. However, the generated speech still fails to generate speech akin to spontaneous human speech in the real world [4, 6].
One of the significant reasons for this limitation is that current speech generation models are trained on speech datasets which have their root in audiobooks [7, 8]. Those datasets typically are characterized by formal reading styles. However, speech from real humans, especially in casual or conversational contexts, rarely adheres to such standardized patterns. Instead, it exhibits more diverse and spontaneous speaking styles, including breathing, pausing, repetitions, changes in speed, and varying emotions. Consequently, there is a pressing need for a new dataset that encompasses more diverse speech styles to advance the field towards generating more spontaneous and human-like speech.
However, directly using in-the-wild speech data is not feasible due to variations in length and quality, frequent background noise, music, reverberation, the presence of multiple speakers within a single sample, and the lack of necessary annotations such as text transcriptions [9]. Training with such data may degrade the performance of speech generation models. While previous works [9, 10] propose automatic preprocessing pipelines to address these issues, they rely heavily on proprietary models, making their pipelines less accessible to the broader community. Additionally, the processing speed of these pipelines remains unknown. An ideal preprocessing pipeline for in-the-wild speech data should process quickly to handle large amounts of data efficiently, allowing for significant dataset scaling. Furthermore, the resulting datasets from these pipelines are limited to monolingual (Chinese-only) data and are relatively small in size (39 hours for [9], 12k hours for [10]).
In response to these issues, we present Emilia-Pipe, the first open-source preprocessing pipeline, which consists of six preprocessing steps: standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering. Emilia-Pipe can effectively transforms in-the-wild speech data into high-quality training data with annotations for speech generation. Additionally, Emilia-Pipe incorporates numerous engineering know-hows to improve robustness and efficiency. The resulting pipeline can process 2.50 hours of raw speech data in one minute using an independent server with eight NVIDIA RTX 4090 GPUs. It is also compatible with different languages.
Leveraging Emilia-Pipe, we construct the first multilingual speech generation dataset from in-the-wild speech data, Emilia. Table 1 compares Emilia with several existing speech generation datasets. The key advantages of the Emilia dataset are summarized as follows:
  • Extensive and Multilingual. The Emilia dataset contains over 101k hours of speech data at 24 kHz and covers six languages: English (En), Chinese (Zh), German (De), French (Fr), Japanese (Ja), and Korean (Ko). To the best of our knowledge, it is the largest academic speech generation dataset.
  • Diverse. Unlike previous datasets, the Emilia dataset comprises mostly spontaneous speech, covering a wide range of speaking styles. This diversity is crucial for training high-quality models to generate spontaneous and human-like speech generation models.
  • Dynamic. The Emilia dataset features an automatic and efficient processing pipeline, allowing it to be easily extended in total duration and language coverage by adding user-specified source audios. To our best, this is the first dynamic speech dataset with an open-source data preprocessing pipeline, facilitating large-scale speech generation research for the community.
To validate the effectiveness of Emilia, we train two text-to-speech (TTS) models on the English subset of the Emilia dataset and compared them with their counterparts trained on the Multilingual LibriSpeech (MLS), a high quality dataset derived from audiobooks. Experimental results from both subjective and objective evaluations demonstrate that Emilia is effective for training high-quality, spontaneous, and human-like speech generation models. Additionally, models trained with the full Emilia dataset show promising performances in multilingual TTS.
The Emilia-Pipe and Emilia dataset, is now publicly available at https://github.com/open-mmlab/Amphion/tree/main/preprocessors/Emilia.
Table 1: A comparison of Emilia with existing datasets for speech generation. Note that the pipeline in [9] and [10] is not publicly-available.
Dataset Data Source Total Duration (hours) Lang. Samp. Rate (Hz) Dynamic
LJSpeech [11] Audiobook 24 En 22.05k
AutoPrepWild [9] In-the-wild 39 Zh 24k/44.1k ✓(not open-source)
VCTK [12] Studio Recording 44 En 48k
Aishell-3 [13] Studio Recording 85 Zh 44.1k
LibriTTS [14] Audiobook 585 En 24k
GigaSpeech [15] In-the-wild 10k En 16k
WenetSpeech4TTS [10] In-the-wild 12k Zh 16k ✓(not open-source)
MLS [8] Audiobook 51k En/Fr/De/Nl/Es/It/Pt/Pl 16k
Libri-Light [7] Audiobook 60k En 16k
Emilia In-the-wild 101k En/Zh/De/Fr/Ja/Ko 24k
2 The Emilia-Pipe Processing Pipeline
This section details Emilia’s data processing pipeline, Emilia-Pipe. As illustrated in Fig. 1, Emilia-Pipe includes six steps, i.e., Standardization, Source Separation, Speaker Diarization, Fine-grained Segmentation by VAD, ASR, and Filtering.
Fig. 1: An overview of the Emilia-Pipe processing pipeline.
2.1 Standardization
The raw speech data in the wild often vary in encoding formats, sampling rates, etc. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16-bit and adjust the target decibels relative to full scale to -20 dBFS. The actual gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. These steps ensure a consistent data format for further processing.
2.2 Source Separation
The raw speech data in the wild often contain background music, which negatively impacts the performance of speech generation models [16, 17]. To address this, we use the source separation technique to extract clean human vocals. Specifically, we utilize the open-source library, Ultimate Vocal Remover, [^1] https://github.com/Anjok07/ultimatevocalremovergui and its pre-trained model, UVR-MDX-Net Inst 3 [18]. [^2] https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models) This model achieves a high signal-to-distortion ratio of 11.15 for vocal separation on the Synth MVSep dataset. Using this model, we effectively separate human vocals from raw speech data for further processing.
2.3 Speaker Diarization
After extracting clean human vocals from the raw speech data, we apply the speaker diarization technique to partition the long-form speech data into multiple utterances based on the speaker. This process generates a series of utterances for each speech data, with each utterance containing only one speaker, ensuring compatibility with existing datasets for speech generation [7, 8, 12, 14]. To achieve this, we leverage the “pyannote/speaker-diarization-3.1” pipeline. [^3] https://github.com/pyannote/pyannote-audio This pipeline includes three core components: speaker segmentation, speaker embedding, and clustering, achieving state-of-the-art speaker diarization performance [19, 20]. The output of this pipeline is a list of temporal annotations indicating the start and end times of the single-speaker utterances.
2.4 Fine-grained Segmentation by VAD
Although the speaker diarization pipeline provides a coarse segmentation for the raw speech data, the resulting utterances may still be too long to fit into memory. To address this, we use a voice activity detection (VAD) model to further segment the utterances into smaller segments ranging from 3 to 30 seconds. This is achieved by concatenating consecutive chunks containing voice activity from the same speaker. We leverage the open-source library Silero-VAD. [^4] https://github.com/snakers4/silero-vad The pre-trained model provided in Silero-VAD achieves a ROC-AUC score of 0.99 on the LibriParty dataset, ensuring accurate detection of voice activity.
2.5 ASR
The absence of reliable text transcriptions impedes the direct use of the Emilia dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [21], which is a state-of-the-art multilingual ASR model capable of speech translation and language identification. To further enhance efficiency, we leverage the WhisperX [22], which builds on faster-whisper [^5] https://github.com/SYSTRAN/faster-whisper backend and the CTranslate2666https://github.com/OpenNMT/CTranslate2 inference engine, is up to four times faster than the official Whisper implementation for the same accuracy while using less memory. Additionally, we omit the original model’s VAD component by using the results from Sec. 2.4 to avoid redundant processing. We also develop batched inference for the faster-whisper backend to transcribe the speech data in parallel. These improvements, combined with the powerful Whisper model, allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.
2.6 Filtering
In real-world scenarios, some noise may not be effectively handled by source separation, the ASR step may introduce errors, and some raw speech data may be of low quality [9]. Therefore, to ensure the quality of the resulting dataset, we filter the data using the following criteria. [^7] Please note that the filtering criteria can be adjusted to fit the specific needs of different use cases. Firstly, we utilize the language identification results from the Whisper Model in Sec. 2.5. We discard the speech data that are not predicted to belong to our target languages (English, French, German, Chinese, Japanese, Korean) and have model language confidence lower than 80%. Secondly, we use DNSMOS P.835 OVRL score [23] to estimate the overall speech quality, preserving only those speech data with the score higher than 3.0. Finally, for each raw speech data, we compute the average character duration over its corresponding segments. We consider segments with an average phone duration outside 1.5 times the interquartile range (IQR) above the third quartile or below the first quartile as outliers and discard the speech data for these segments. After filtering, we obtain the resulting dataset, which are now ready for training the speech generation model.
2.7 Performance Evaluation
Table 2: Statistics of 600 hours in-the-wild speech data processed by Emilia-Pipe.
Dataset Duration (s) DNSMOS P.835 OVRL Total Duration (hours)
min max avg ± std min max avg ± std
Raw 9.22 18,056.98 1,572.53 ± 1,966.66 1.08 3.51 2.50 ± 0.62 598.87 (100.00%)
Processed w/o Filtering 1.00 30.00 7.18 ± 5.06 0.91 3.67 2.86 ± 0.51 340.54 (56.86%)
Processed 3.00 30.00 8.98 ± 4.99 3.00 3.67 3.26 ± 0.14 176.22 (29.43%)
To analyze Emilia-Pipe, we randomly sample a subset of raw speech data, approximately 600 hours, and use Emilia-Pipe to process this subset to evaluate its effectiveness and efficiency.
The evaluation is conducted on an independent server with eight NVIDIA RTX 4090 GPUs. The whole processing time takes about 3.99 hours. Table 2 shows the processing results of Emilia-Pipe on this subset. The raw data has a wide range of audio durations from 9.22 to 18,056.98 seconds, with an average of 1,572.53 seconds and high variability. The DNSMOS P.835 OVRL scores range from 1.08 to 3.51, with an average of 2.50, indicating varied overall quality. After filtering, the total duration for the resulting data is further reduced to 176.22 hours, retaining 29.43% of the raw speech data, and the average DNSMOS P.835 OVRL score significantly improves to 3.26 with minimal variability, indicating that Emilia-Pipe can effectively transform in-the-wild speech data into high-quality training data for speech generation. Besides, for processing this subset, Emilia-Pipe processes about 2.50 hours of data every one minute. This demonstrates that our processing method significantly exceeds real-time standards, making it ideal for preprocessing extensive speech data and scaling up the training dataset.
3 The Emilia Dataset
3.1 Overview
[…]
Fig. 2: Duration statistics (hours) of the speech data by language.
Using Emilia-Pipe, we construct the Emilia dataset from a vast collection of speech data sourced from diverse video platforms and podcasts on the Internet, covering various content categories such as talk shows, interviews, debates, sports commentary, and audiobooks. This variety ensures the dataset captures a wide array of real human speaking styles.
After processing, the initial version of the Emilia dataset includes a total of 101,654 hours of multilingual speech data in six different languages: English, French, German, Chinese, Japanese, and Korean. Fig 2 provides the duration statistics for each language in the dataset.
3.2 Dataset Analysis
In this subsection, we analyze the quality and diversity of the Emilia.
3.2.1 Quality
To evaluate quality, we compared Emilia with existing datasets using DNSMOS P.835 OVRL scores. This non-intrusive speech quality metric reflects the overall quality of the speech data and is highly correlated with human ratings [23]. Table 3 presents the speech quality comparison between Emilia and several existing datasets.
Emilia achieves a DNSMOS P.835 OVRL score of 3.26, ranking third among all datasets. The results indicate that, despite being sourced from raw speech data in the wild, after preprocessing, the speech quality of the Emilia dataset is comparable to existing datasets sourced from studio recordings or audiobooks and outperforms all existing datasets sourced from in-the-wild speech data.
Table 3: Quality comparison between Emilia and nine existing datasets. The scores for LJSpeech, AutoPrepWild, Aishell-3, and LibriTTS are derived from [9]. The score for Libri-Light is computed from its official "small" subset, and the score for WenetSpeech4TTS is computed from its official "basic" subset. The scores for MLS and Emilia are computed from a randomly sampled 600-hour subset.
Dataset DNSMOS P.835 OVRL
LJSpeech [11] 3.30 ± 0.17
AutoPrepWild [9] 3.24 ± 0.21
VCTK [12] 3.20 ± 0.18
Aishell-3 [13] 3.15 ± 0.17
LibriTTS [14] 3.25 ± 0.19
GigaSpeech [15] 2.52 ± 0.54
WenetSpeech4TTS [10] 3.18 ± 0.22
MLS [8] 3.33 ± 0.19
Libri-Light [7] 3.25 ± 0.26
Emilia 3.26 ± 0.14
3.2.2 Diversity
The Emilia dataset comprises a collection of speech data from a wide range of video platforms and podcasts, capturing diverse speaking styles of real human speech. To quantify this diversity, we conducted analyses on both the acoustic and semantic feature space, comparing it with the MLS dataset, which is derived from audiobooks and widely used for training speech generation models.
Fig. 3: A comparison of acoustic and semantic diversities between Emilia and MLS datasets.
Specifically, we randomly select 5,000 samples each from the English subset of MLS and Emilia. To analyze the diversity of acoustic features, we leverage a pre-trained WavLM model [^8] https://huggingface.co/microsoft/wavlm-base-plus to extract acoustic representations, capturing a variety of acoustic characteristics such as speaker, emotion, and prosody [24]. We then apply the PCA algorithm to reduce the dimensionality of these representations to two. As shown in Fig. 3 (a), the Emilia dataset exhibits a broader dispersion, contrasting with MLS, which shows a more compact clustering. The more scattered pattern highlights the Emilia dataset as encompassing a richer acoustic characteristic coverage compared to the MLS dataset derived from audiobooks.
For the semantic diversity analysis, we employ a pre-trained Sentence-BERT model [^9] https://github.com/UKPLab/sentence-transformers to generate text representations for the transcripts of each speech data. Consequently, each speech data is represented as a 768-dimensional vector based on its textual content, providing a comprehensive approximation of its semantic. Similar to the analysis above, we reduced the dimension of the semantic features to two. As shown in Fig. 3 (b), the scatter of textual features indicates that the Emilia dataset covers a wide range of textual content, validating the significant diversity in Emilia’s semantic coverage.
Table 4: Objective and subjective evaluation of TTS models using Emilia and MLS on LibriSpeech-Test and Emilia-Test evaluation sets.
Model Train Set LibriSpeech-Test Emilia-Test
WER $\downarrow$ SIM-O $\uparrow$ FSD $\downarrow$ CMOS $\uparrow$ SMOS $\uparrow$ WER $\downarrow$ SIM-O $\uparrow$ FSD $\downarrow$ CMOS $\uparrow$ SMOS $\uparrow$
SoundStorm MLS 8.9% 0.612 49.11 0.360.36-0.36- 0.36 $3.133.13$ 7.7% 0.587 20.76 $0.090.09$ $3.713.71$
Emilia 8.4% 0.577 24.73 0.190.19-0.19- 0.19 $3.283.28$ 6.6% 0.618 12.73 $0.190.19$ $3.733.73$
VoiceBox MLS 6.1% 0.625 16.83 $0.360.36$ $3.623.62$ 8.2% 0.528 15.94 $0.280.28$ $3.613.61$
Emilia 7.2% 0.585 23.24 $0.420.42$ $3.773.77$ 7.4% 0.601 14.07 $0.280.28$ $3.763.76$
4 Experiments
In this section, we evaluate the effectiveness of the Emilia dataset in TTS applications through two experiments: the English-only experiment and the multilingual experiment. In the English-only experiment, we compare the performance of TTS models trained with the English subset of the Emilia dataset to those trained with the English subset of the MLS dataset. In the multilingual experiment, we train the models with the full Emilia dataset, which comprises 101,654 hours of multilingual speech data, and evaluate their multilingual TTS performance.
4.1 Experimental Setups
4.1.1 Baselines
In the experiments, we implement two TTS models as baselines: SoundStorm [2] and VoiceBox [16].
SoundStorm uses text to predict speech semantic tokens [25] in an autoregressive manner, and then generates acoustic tokens from a neural audio codec with bidirectional attention and confidence-based parallel decoding [2]. VoiceBox is a non-autoregressive speech synthesis model that adopts the non-autoregressive flow-matching method and the transformer for speech generation tasks. This model efficiently generates speech by learning the distribution of the mel-spectrogram conditioned on both text input and speech context.
4.1.2 Evaluation Metrics
To evaluate the baselines, we conduct both objective and subjective evaluations.
For the objective evaluation, we consider the following aspects: (1) Intelligibility: Measured by the Word Error Rate (WER) of the synthesized speech’s transcription compared to the input text. For LibriSpeech-Test, we use a finetuned HuBERT-Large ASR model. [^10] https://huggingface.co/facebook/hubert-large-ls960-ft For other testsets, we use the Whisper-medium model. [^11] https://huggingface.co/openai/whisper-medium (2) Coherence: Assessed by speaker similarity between generated speech and the speech prompt using the WavLM-TDCNN speaker embedding model. [^12] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification We report similarity to the original speech prompt (SIM-O). (3) Naturalness: Evaluated using the Fréchet Speech Distance (FSD), which measures the similarity between the distributions of generated and real samples in a feature space. A lower FSD indicates higher speech quality and diversity [16]. We adapt the metric for speech by using the emotion2vec features. [^13] https://github.com/ddlBoJack/emotion2vec/tree/main
For the subjective evaluation, we randomly select eight samples each from LibriSpeech-Test and Emilia-Test. Twelve proficient English speakers served as judges. we use SMOS (Similarity Mean Opinion Score) to evaluate the speaker similarity of the speech to the original speech prompt. The SMOS scale ranges from 1 to 5, with increments of 0.5 points. CMOS (Comparative Mean Opinion Score) is used to evaluate the comparative naturalness of the synthesized speech against a given speech prompt. The CMOS scale ranges from -3 (indicating the synthesized speech is much worse than the speech prompt) to 3 (indicating the synthesized speech is much better than the speech prompt), with intervals of 1.
4.2 Emilia English versus MLS English
The experiment evaluates the effectiveness of the proposed Emilia dataset by comparing the performance of models trained on the English subsets of both Emilia and the MLS dataset, a high-quality dataset derived from audiobooks. The total duration of the Emilia dataset is 46,000 hours, while the MLS dataset comprises 44,500 hours. The size of the datasets can be considered roughly equivalent. To thoroughly evaluate the models, we utilize the LibriSpeech-Test evaluation set, containing 1,200 speech samples in formal reading styles akin to those of MLS, and the Emilia-Test evaluation set, which includes 600 speech samples in diverse spontaneous speaking styles. Both test sets contain speech data that are unseen by the baselines.
Table 4 presents the results of both objective and subjective evaluations for the Emilia and MLS datasets on the LibriSpeech-Test and Emilia-Test. From these results, we observe that models trained on both the Emilia and MLS datasets demonstrate similar levels of speaker similarity (measured by SIM-O and SMOS) and intelligibility (measured by WER). This suggests that the Emilia dataset, despite being sourced from raw speech data in the wild, is as effective as high-quality datasets derived from audiobooks after processing with our proposed Emilia-Pipe. The VoiceBox model trained with both datasets demonstrates a similar level of naturalness. In contrast, the SoundStorm model shows significant improvement in FSD and CMOS on the Emilia-Test, which contains speech prompts in diverse spontaneous speaking styles. These results may indicate that autoregressive TTS models benefit more from speech with diverse speaking styles compared to non-autoregressive models.
4.3 Multilingual TTS
Table 5: Objective evaluation of TTS models trained on Emilia on six languages.
Language Model WER $\downarrow$ SIM-O $\uparrow$ FSD $\downarrow$
En SoundStorm 6.2% 0.614 14.82
VoiceBox 5.8% 0.581 14.47
Zh SoundStorm 4.1% 0.564 35.59
VoiceBox 4.7% 0.560 48.95
De SoundStorm 6.8% 0.680 32.72
VoiceBox 13.3% 0.633 44.68
Fr SoundStorm 8.2% 0.589 42.92
VoiceBox 17.5% 0.550 48.82
Ja SoundStorm 3.6% 0.625 49.42
VoiceBox 10.8% 0.562 49.39
Ko SoundStorm 10.9% 0.681 47.93
VoiceBox 15.2% 0.608 63.00
Next, we conduct experiments to evaluate the performance of the multilingual TTS capability of the baselines trained on the full Emilia datasets, encompassing six languages. The test set for En is Emilia-Test. The test set for Zh is from Aishell-3. The test sets for De, Fr, Ja, and Ko are from Common Voice. Each test set contains at least 500 samples. Due to space limitations, we only present the objective evaluation results in Table 5, omitting the subjective evaluation results. The findings confirm that both models exhibit strong zero-shot multilingual TTS performance, underscoring the multilingual effectiveness of the Emilia dataset.
5 Conclusions
In conclusion, this paper introduces Emilia, an extensive, multilingual, and diverse dataset for speech generation, along with Emilia-Pipe, an open-source preprocessing pipeline that can effectively and efficiently transforms raw speech into high-quality training data, facilitating the scale-up of the Emilia dataset. The initial version of the Emilia dataset includes over 101k hours of speech data in six languages, featuring a wide variety of in-the-wild real-world speech. Both objective and subjective evaluations confirm the effectiveness of the Emilia dataset. This work aims to advance speech generation towards producing high-quality, spontaneous, and human-like speech, and we invite the research community to utilize Emilia-Pipe for large-scale speech generation.